Rafa's age is 6 years less than Rosa's what is the age of each of the sum if their ages is 24
Answer by Cromlix(4381)

Hi there,
Rosa's age = x
Rafa's age = x - 6
Sum of the ages = 24
x + x - 6 = 24
2x = 24 + 6
2x = 30
x = 30/2
x = 15
Rosa's age = x = 15
Rafa's age = x-6 = 15 - 6 = 9
Rasa's age = 15
Rafa's age = 9
Hope this helps:-)